greenhab report – november 12th

Crew 198 GreenHab Report – 12-NOV-2018

GreenHab Officer: Andrei

Environmental control: (Choose which of the following is appropriate and explain further if needed)


Shade cloth on

Average temperatures:

Low: 70

High 80

Hours of supplemental light: none

Daily water usage for crops: 4 gallons

Water in Blue Tank—(full) unknown number of gallons

Time(s) of watering for crops: 0900

Changes to crops: Note all emergence of seedlings, death of seedlings, etc (These are also mandatory to write in the GreenHab book for 2018-2019) none

Narrative: no other information at this time

Harvest: This is our research so it is important that all harvested crops are noted and also weighed and that weight recorded in the GreenHab book and here.

Support/supplies needed:none

sol summary – november 12th

Crew 198 Sol Summary Report 11-12-2018

Sol: 2

Summary Title: Successful EVA and Team Building

Author’s name: Jayne Paulson

Mission Status: Things are going very well.

Sol Activity Summary: We learned about the role of flight surgeons in aerospace medicine, ultrasound in spaceflight, and winter and water survival.

Look Ahead Plan: Tomorrow we are launching a rocket!

Anomalies in work: Generator as described

Weather: Cool in the morning, sunny and perfect for the EVA

Crew Physical Status: Good, no complaints

EVA: Kissing Camel Ridge

Reports to be file: Operations, Green Hab, Journalist, EVA Request, EVA Report, Sol Summary, Crew Bios

Support Requested: None

Operations Report – November 11th 2018

Crew 198 Operations Report 11/10/2018

SOL: 1

Name of person filing report: Dana Levin

Non-nominal systems: Engineering Lock light and power plugs near back section of habitat

Notes on non-nominal systems: they’ve been non functional since our first mission 5 years ago at this point it may be more accurate to refer to the non functional state as “nominal” I will let you know if they suddenly become “off-nominal” and turn on.

Generator (hours run): 7 hours (midnight to 0700)

Solar— SOC 65% (Before generator is run at night)

Diesel – 60%

Propane – psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 15 gallons.

Water (trailer) – 400 gallons.

Water (static) – 60 gallons

Trailer to Static Pump used – yes

Water (loft) – Static to Loft Pump used – yes

Water Meter:

ATV’s Used: (Honda, 350.1, 350.2)

Oil Added? No

ATV Fuel Used: 0.5 Gals each

# Hours the ATVs were Used today: .75h

Notes on ATVs: HOnda ATV will not shift into reverse

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: disabled

Hours: 0.75

Beginning charge: 100

Ending charge: 90

Currently charging: yes

Spirit rover used: yes

Hours: 0.75

Beginning charge: 100

Ending charge: 89

Currently charging: yes

Opportunity rover used: yes

Hours: 0.75

Beginning charge: 100

Ending charge: 88

Currently charging: yes

Curiosity rover used: yes

HabCar used and why, where? Yes to pick up water from hollow mountain (2 runs)

General notes and comments:

Summary of internet: functional

Summary of suits and radios: suits functional, radios not used/ 2MARS crews use our own comms

Summary of Hab operations: today was orientation and medical scenarios

Summary of GreenHab operations: maintenance

Summary of ScienceDome operations: maintenance and engineering

Summary of RAMM operations: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: none at present.

Sol Summary Report – November 11th 2018

Crew 198 Sol Summary Report 11Nov2018

Sol 1

Summary Title The day was very educational with didactic sessions supported by real time implementation of skills learned.

Author’s name Jayne Paulson

Mission Status: Team morale is high and they seem to be bonding well. They are working well as a team and identified appropriate roles within the team.

Sol Activity Summary: am didactic sessions and afternoon EVA

Look Ahead Plan: Our debrief today identified significant strengths and areas for improvement on future EVAs.We plan to implement these changes on EVA tomorrow.

Anomalies in work: none

Weather:Partly cloudy

Crew Physical Status: Crew physical status is excellent, and everyone is excited to be on Mars!

EVA: Recovery of Chinese Satellite

Reports to be filed:

Sol Summary

Engineering Report

Journalist Report (Photos)

GreenHab Report

EVA Request

Crew Bios

Support Requested: None

GreenHab Report – November 11th 2018

Dana Levin


Environmental control:

Ambient (no heating or/cooling)

Shade cloth off

Average temperature: (N/A until new sensor is delivered)

Changes to crops: Crops appear to be alive, no unexpected plants, animals, or other life has appeared

Daily water usage for crops: 4 gallons

Time(s) of watering for crops: 1730

Moringa research observations: (Currently not operational)

Changes to research plants: none

Daily watering and amount of water used:

Aquaponics: (Currently not operational)

Narrative: Today was the orientation to the greenhab, we now know how to take care of the plants and look forward to doing it in the future

Support/supplies needed: more water

Copyright © The Mars Society. All rights reserved. | Main Site